# Phase field crystal modelling of deformation assisted precipitation in binary alloys

### v1

@@ Abstract
A third mechanism observed at high $\dot{\epsilon}$ deformations is recrystallization at the grain boundary, ie. the formation of lower-strain A-rich grains. Similar behaviour has recently been modelled in a pure material phase-field crystal model under laser deposition [32]. A typical view of recrystallization events we observe are shown in the orientation field of Figure 2. Due to working in two spatial dimensions these recrystallized regions are not long lived. Indeed, upon performing identical simulations in the scenario where solute is frozen ($M_{c},\xi_{c}=0$), we see the recrystallized regions initially attain an orientation intermediate to the surrounding bulk grains and eventually coarsen, post nucleation. The mechanism of formation for these recrystallized grains is achieved through collective motion of nucleated dislocations to buckled regions of the grain boundary. In the absence of precipitation, this mechanism has been observed experimentally in the nucleation of twins during continuous deformation of copper bicrystals [33].
@@
To summarize, the collective action of the above three phenomena causes $f$ to exhibit a plateau-like behaviour. As can be seen in Figures 2 and 5, while not halted completely, the maximum in $df/dt$ discontinuously transitions from one to a pair of discrete maxima. Such behaviour in Avrami curves is commonly associated with multiple transformation mechanisms of similar driving forces competing with one another. Our findings have been observed experimentally in the situation when recrystallization is halted by precipitation [34] and in the precipitate fraction of an Al-Zn-Mg-Cu during hot deformation [3]. It is noted, however, that strain-rate dependence in the present study cannot be quantitatively compared to the mechanisms proposed in [3] as defects cannot frustrate in two dimensions.
@@
It is noteworthy that the three mechanisms described above induce an overall acceleration of the precipitation reaction as it becomes progressively easier to partition the solute in the bulk toward its equilibrium values. As shown in the data of Figure 4, precipitation under large deformation occurs over a shorter time-scale than the case of low (or no) deformation rate below some critical value. These results are also consistent with the typically observed experimental behaviour of TTT diagrams in the presence of deformation [1, 35].
To conclude, we have shown that modified mobilities may be introduced into an improved XPFC alloy model such as those in Eq. (6) to provide a robust atomistic-continuum framework for examination of first order phase transformations in the solid state. We have further shown that upon application of deformation to a bicrystal, beyond an initial stage of elastic compensation where the precipitation largely occurs along a flat grain boundary, new defect assisted nucleation mechanisms begin to compete with classic precipitation mechanisms along the grain boundary. Our results provide an atomistic perspective to theß phenomenon of accelerated precipitation commonly observed in deformed systems. We also observe recrystallization events along the grain boundary, which coarsen due to the lack of defect frustration, which is not expected to occur in three dimensions.
I Acknowledgements

### v2

@@ IV Results & Discussion
A third mechanism observed at high $\dot{\epsilon}$ is recrystallization at the GB, ie. the formation of lower-strain A-rich grains. These grains further serve to lower the barrier for precipitate nucleation. Similar behaviour has recently been modelled in a pure material PFC model under laser deposition [32, 33]. A typical view of recrystallization events we observe are shown in the orientation field of Figure 2. Due to working in two spatial dimensions these recrystallized regions are not long lived. Indeed, upon performing identical simulations in the scenario where solute is frozen ($M_{c},\xi_{c}=0$), we see the recrystallized regions initially attain an orientation intermediate to the surrounding bulk grains and eventually coarsen, post nucleation. The mechanism of formation for these recrystallized grains is achieved through collective motion of nucleated dislocations to serrated regions of the GB. In the absence of precipitation, this mechanism has been observed experimentally in the nucleation of twins during continuous deformation of copper bicrystals [34].
@@
To summarize, the collective action of the above three phenomena causes $f$ to exhibit a plateau-like behaviour. As shown in Figures 2 and 5, the maximum in $df/dt$ discontinuously transitions from one to a pair of discrete maxima. Such behaviour in Avrami curves is commonly associated with multiple transformation mechanisms of similar driving forces competing with one another.
Our findings have been observed experimentally in the situation when recrystallization is halted by precipitation [35] and in the precipitate fraction of an Al-Zn-Mg-Cu during hot deformation [3]. These experimental results have also been examined from the perspective of a Kampmann-Wagner model equipped with a vacancy concentration and an evolving dislocation density [36]. However, their results do not consider nucleation of precipitates on dislocations. Experimentally, such dislocations are expected to be released during simultaneous deformation and precipitation [37]. It is noted, however, that strain-rate dependence in the present study cannot be quantitatively compared to the mechanisms proposed in [3] as defects cannot frustrate in two dimensions. Nonetheless, our results suggest that an additional strain-induced dislocation nucleation mechanism may be required to fully model the dynamics of precipitation.
@@
Our findings also indicate that the continuous application of a deformation may induce an acceleration of the precipitation reaction. This is because the introduction of defects enhances precipitate nucleation in the bulk. As shown in the data of Figure 4, precipitation under large deformation occurs over a shorter time-scale than the case of low (or no) deformation rate, beyond which we observe a bifurcation in $df/dt$. These results are also consistent with typically observed experimental behaviour of TTT diagrams in the presence of deformation [1, 38].
V Conclusion
To conclude, we have shown that modified mobilities introduced into an improved XPFC alloy model as in Eq. (6) provide a robust atomistic-continuum framework for examination of first order phase transformations in the solid state. This framework allows for temperature based control of the activity of structural vs. diffusively driven processes that may be tuned independently to driving forces arising from the free energy.
We have further shown that upon application of deformation to a bicrystal, beyond an initial stage of elastic compensation where precipitation occurs along a flat GB, new defect assisted nucleation mechanisms begin to compete with classic precipitation mechanisms along GBs. Our results demonstrate that these mechanisms alter the dispersion of precipitates within the system and introduce a second time-scale into the Avrami dynamics of these systems.
The results of this study demonstrate how PFC modeling may be used to elucidate how precipitate matrix interactions on multiple time and length scales affect microstructure evolution. In regards to hot deformation, information about atomistic properties such as the dislocation density and GB behaviour can be further investigated. Additionally, future studies in three spatial dimensions may be able to shed further light on both recrystallization and defect pileups in such systems.
Acknowledgements.
@@
Appendix A Free Energy Parametrization
The parameters we employ that specify the ideal, mixing and gradient components of the free energy in Equation 1 to our particular material of interest are shown in Table 1. The parameters we employ for the correlation function in Equation 3 are shown in Table 2. These parameters are held fixed over the course of every simulation we perform. The respective phase diagram corresponding to these parameters is also shown in Figure 6. Phase diagram construction is performed by numerically identifying common tangents in the free energy landscape after an assumed mode expansion for the density is inserted into the free energy and a subsequent minimization of the mode expansion amplitudes is performed as described in c.f. [16, 15].
$c_{0}$ $\eta$ $\chi$ $\omega$ $W_{c}$
0.5 1.4 1 0.005 0.25
Table 1: Free energy parameters employed in Eq. (1)
$k_{A}$ $k_{B}$ $T_{A}^{0}$ $T_{B}^{0}$ $\sigma_{A,0}$ $\sigma_{B,0}$ $\sigma_{A,1}$ $\sigma_{B,1}$ $A_{A,0}$ $A_{B,0}$
[…]
Table 2: Parameters employed for the Fourier space two-point density-density correlation function in Eq. (3)
Figure 6: Phase diagram associated with free energy parameters shown in Table 1 with respect to dimensionless temperature, $T/T^{0}$ and concentration $c$. The average density is fixed here at 0.05 which is sufficiently distant from any phases outside of the liquid and solid phase.
Appendix B Numerical Implementation
Dynamical evolution of Eqs. 4 and 5 in the main text was performed with a semi-implicit Fourier spectral method [15]. The time discretized form we adopt for our equations in Fourier space is,
$\frac{\hat{n}^{t+1}-2\hat{n}^{t}+\hat{n}^{t-1}}{\Delta t^{2}}+\beta\frac{\hat{n}^{t+1}-\hat{n}^{t}}{\Delta t}$ $=\hat{L}_{n}^{t}\hat{n}^{t+1}-k^{2}\hat{N}_{n}^{t},$
$\frac{\hat{c}^{t+1}-\hat{c}^{t}}{\Delta t}$ $=\hat{L}_{c}^{t}\hat{c}^{t+1}-k^{2}\hat{N}_{c}^{t}.$ (8)
Here, $\hat{L}_{n/c}^{t}$ and $\hat{N}_{n/c}^{t}$ are the linear and non-linear terms associated with each field in Fourier space and superscripts indicate the time-step of each term. In the above, we apply the frequently applied approximation that the linear operator associated with the evolution of the $n$ and $c$ fields is diagonal with elements $\hat{L}^{t}_{n/c}$. Rearranging the above, we can rewrite the discretized scheme as follows,
$\hat{n}^{t+1}$ $=\frac{\hat{n}^{t}(2+\beta\Delta t)-k^{2}\hat{N}_{n}^{t}\Delta t^{2}-\hat{n}^{t-1}}{(1+\beta\Delta t-\hat{L}_{n}^{t}\Delta t^{2})}$ (9)
$\hat{c}^{t+1}$ $=\frac{\hat{c}^{t}-k^{2}\hat{N}_{c}^{t}\Delta t}{(1-\hat{L}_{n}^{t}\Delta t)}.$ (10)
The fields can then be evolved using the information from the previous time-step for $c$ and previous two time-steps for $n$. The parameters employed to evolve the dynamical equations are shown in Table 3. These parameters were employed to allow for simultaneous precipitation and plastic response for the strain-rates we employ. The choice of strain-rate was set such that a non-negligible amount of strain-energy could be imparted into the system prior to precipitation while staying within the confines of strains applied during deformation of materials.
Each choice of temperature, strain-rate and activation energy were simulated using the same discretization scheme. We employ a two-dimensional system box of $N_{x}\times N_{y}$ pixels with a discretization of $dx$ and $dy$ along each respective axis. Time was discretized with an increment $dt$. The parameters we employ for our simulations are shown in Table 4.
The initial discretizations ($dx^{0}$, $dy^{0}$) along each axis are varied along with the initial grain boundary misorientation to allow for an integer number of atoms initially present along both axes inside the simulation box. These variations are kept within the ranges of $12.00^{\circ}\pm 0.05^{\circ}$ and $0.10\pm 0.01$ for the misorientation and discretization respectively. We do not observe significant deviations in the behaviour precipitated fraction or the pressure over different parameter choices within our set tolerance (solutions are typically not unique within the tolerance range).
Application of deformation is performed numerically through periodic variation of the spatial discretization while preserving the overall volume of the system. The discretization is varied according to the following scheme,
$dx^{t+T}$ $=dx^{t}(1+\dot{\epsilon}Tdt)$ (11)
$dy^{t+T}$ $=dy^{t}/(1+\dot{\epsilon}Tdt)$ (12)
where T denotes the number of numerical time-steps between each application of deformation.
$M_{n}^{0}$ $M_{c}^{0}$ $E_{A}^{c}$ $E_{A}^{n}$ $\beta$ $N_{n}$ $N_{c}$
0.02 1 0.1-1 10 0.1 0.05 0.01
Table 3: Parameters employed for dynamical evolution of Eqs. 4 and 5.
$N_{x}$ $N_{y}$ $dx$ $dy$ $dt$ $T$
4096 4096 $0.1\pm 0.01$ $0.1\pm 0.01$ 0.5 300
Table 4: Numerical system size and discretization employed along the x and y axes. The initial grain boundary misorientation and discretizations vary slightly between simulations to accommodate an integer number of atoms parallele to the GB for varying supersaturation.